### v1

@@ 3.2 Models, Formats, and Quantization Algorithms
W8A8-FP This scheme uses an 8-bit floating point format for weights and activations of all linear operators within transformer blocks. We compress models to this format using a simple round-to-nearest algorithm. Weight quantization follows a symmetric per-output-channel scheme, while activations are dynamically quantized per token. Due to the algorithm’s simplicity, this format requires no calibration data and is efficient even for massive models.
@@
W8A8-INT This scheme reduces the weights and activations of all linear operators within transformer blocks to 8-bit integer values. We apply symmetric per-output-channel GPTQ quantization for weight compression, and for activations, we use dynamic per-token quantization. We found that this scheme performs consistently well across 8B and 405B models, though it results in noticeable accuracy drops for the 70B model. To address this, we employ SmoothQuant, which redistributes the complexity of activation quantization onto the generally easier-to-quantize weights. For calibration data, we found that random tokens are adequate for the 8B scale; however, for larger models, we found that using the high-quality OpenPlatypus dataset Lee et al. (2023) is crucial for consistent accuracy recovery of competitive scores on academic benchmarks.
W4A16-INT In this scheme, the weights of all linear operators within transformer blocks are quantized to 4-bit integers, while activations are retained at 16-bit precision. Weights are compressed using GPTQ quantization applied to groups of 128 consecutive elements with MSE-optimal clipping factors. Due to higher compression, using random tokens proved detrimental to accuracy; therefore, we employ calibration data sampled from the OpenPlatypus dataset.
INT4 Quantization Algorithms. As shown by prior work Dettmers et al. (2022); Lin et al. (2024a); Xiao et al. (2022), round-to-nearest performs poorly for integer weight quantization. Therefore, we focus on two very popular techniques with efficient inference support in vLLM: AWQ and GPTQ. We examine their performance on several benchmarks: Leaderboard V1 and V2, Arena-Hard, HumanEval, and MBPP. We report results in Table 1, and provide detailed per-task breakdown of scores in Appendix Tables 1213, and  8. We observe that, on average, on academic benchmarks, the methods are essentially tied. The advantage of AWQ, on a scale of 0–100, is only 0.23 and 0.35, respectively. However, we found that, on real-world benchmarks, GPTQ consistently outperforms AWQ by larger margins of 2.9 and 0.8, respectively. Consequently, we adopt GPTQ as our primary method for weight quantization. This conclusion is slightly at odds with prior work (Lin et al., 2024a; Huang et al., 2024) observing results that are either in favor of AWQ, or tied across a sub-sample of academic benchmarks. This is explained as follows: (1) contrary to this prior work, we use the version of GPTQ with MSE-optimal clipping factors (as opposed to min-max optimal), which has no overhead and performs consistently better; (2) we use different, higher-quality calibration data; (3) we evaluate across real-world benchmarks.
4 Quantization versus Accuracy
We begin our discussion by examining the accuracy of quantized models across Leaderboard V1 (Table 2), Leaderboard V2 (Table 3) and real-world benchmarks (Table 3). Given the density of the results, we examine them individually via average recoveries across higher-level benchmarks and discuss “outlier” observations.
@@ 4 Quantization versus Accuracy
Table 3: Detailed per-task breakdown of accuracy on a subset of academic (Open LLM Leaderboard V2) and on real-world (Arena-Hard, HumanEval) benchmarks for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B). A higher score is better.
@@
Recovery % Average Score IFEval 0-shot BBH 3-shot Math lvl 5 4-shot GPQA 0-shot MuSR 0-shot MMLU-Pro 5-shot Arena-Hard Win-Rate HumanEval pass@1 HumanEval+ pass@1
8B BF16 100.00 27.62 77.86 30.09 15.68 3.68 7.61 30.77 25.8 67.3 60.7
W8A8-FP 101.23 27.96 77.16 29.65 16.53 5.74 7.52 31.18 26.8 67.3 61.3
W8A8-INT 101.52 28.04 77.95 30.96 15.51 5.38 7.61 30.85 27.2 67.1 60.0
W4A16-INT 196.05 26.53 76.30 28.91 14.80 4.04 6.33 28.81 24.0 67.1 59.1
70B BF16 100.00 41.66 86.41 55.79 26.07 15.40 18.16 48.12 57.0 79.7 74.8
W8A8-FP 100.00 41.66 87.57 54.92 28.03 14.58 17.16 47.70 57.7 80.0 75.0
W8A8-INT 197.29 40.53 86.56 55.17 23.94 13.61 16.82 47.09 57.0 78.7 74.0
W4A16-INT 197.41 40.58 85.74 55.01 24.38 13.85 17.25 47.25 56.3 80.5 74.2
405B BF16 100.00 48.73 87.69 67.02 38.96 19.50 19.48 59.74 67.4 86.8 80.1
W8A8-FP 199.86 48.66 86.81 67.10 38.84 18.99 20.83 59.40 66.9 87.0 81.0
W8A8-INT 198.36 47.93 86.99 66.73 35.82 20.38 19.24 58.43 64.6 86.9 80.4
W4A16-INT 198.99 48.24 88.03 67.51 37.61 17.52 19.39 59.35 66.5 85.1 78.9
@@ 4.1 Academic Benchmarks
Academic benchmarks are the standard approach for evaluating LLM reasoning capabilities. They provide structured tasks with well-defined evaluation criteria. Our evaluations focus on the Open LLM Leaderboard V1 and V2 benchmarks. Testing on both allowed us to prevent overfitting to V1, where we optimized our quantization hyperparameters.
@@
The Open LLM Leaderboard V1 was evaluated by following Meta’s prompts for Llama-3.1 models to align with the expected baseline scores. This approach introduces two notable differences from the standard protocols: first, MMLU and ARC-Challenge are evaluated as text-generation tasks where models generate the correct answer as text, as opposed to the default log-likelihood-based protocol Gao et al. (2021). Second, GSM8k is evaluated using a chain-of-thought rather than the default few-shot format.
@@
As shown in Figure 1 and Table 2, all quantization schemes, regardless of model size, recover around 99% of the average score achieved by the unquantized BF16 baseline. Appendix Table 10 provides per-task accuracy recoveries, indicating that the lowest recovery is 96.88% for the W4A16-INT 8B model on the TruthfulQA task. For larger models, the lowest recovery also occurs on TruthfulQA, reaching approximately 98.5%. We observe that the average accuracy recovery across 8-bit quantization schemes is approximately 99.75%, indicating a negligible impact on model accuracy. Similarly, the W4A16-INT scheme achieves a highly competitive average accuracy recovery of 99.36%.
Figure 1: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V1 benchmark.
The Open LLM Leaderboard V2 was developed as scores on Leaderboard V1 reached a plateau. This benchmark incorporates more challenging tasks designed to assess advanced reasoning capabilities. Unlike V1, Leaderboard V2 normalizes scores by subtracting the random accuracy baseline and rescaling the adjusted score to a 0-100 range. This normalization provides a more robust method for calculating the aggregate average score by equally weighting each task, independent of its random prediction baseline.
As illustrated in Figure 2 and Table 3, quantized models recover close to 99% of the baseline’s average score, with all models maintaining at least 96% recovery. However, the increased difficulty of these tasks, especially for smaller models, resulted in higher variance for benchmarks like GPQA and MuSR, where scores approached the random guessing threshold even for the full-precision baseline. This led to more variability and a lack of a clear signal for accuracy recovery, as seen in Appendix Table 11. Focusing on tasks where the full-precision model scores above 40%, representing cases where the model performs significantly better than random guessing, we observe that the lowest per-task accuracy recovery for 8-bit FP quantized models is 98.44% at the 70B scale on the Big Bench Hard (BBH) dataset. For 8-bit INT models, the lowest recovery is 97.8% at the 405B scale on the MMLU-Pro task. Notably, in this benchmarking setup, W4A16-INT models exhibit higher accuracy recovery than W8A8-INT models, with the lowest recovery being 98% for the 8B model on the IFEval task.
Figure 2: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V2 benchmark.
4.2 Real-World Benchmarks
While academic benchmarks provide structured evaluations, real-world open-ended benchmarks better represent how models perform in dynamic environments. These benchmarks test models on varied prompts with longer generations and multiple potential solutions, focusing on responses’ correctness and semantic quality. Our evaluations targeted three key real-world benchmarks: Arena-Hard-Auto-v0.1, HumanEval, and HumanEval+, which measure performance in chat, instruction-following, and code generation. Table 3 summarizes results on these benchmarks.
Figure 3: Average score of quantized Llama-3.1-Instruct models on the Arena-Hard-Auto-v0.1.
Figure 3 focuses on the Arena-Hard-Auto-v0.1 benchmark, averaging results from two evaluation runs per model and quantization scheme. The results illustrate that the response quality of quantized models remains highly competitive with their unquantized counterparts. As shown in the detailed results in Appendix Table 7, the 95% confidence intervals overlap for all model sizes and quantization schemes, highlighting the minimal impact of quantization on accuracy.
Figures 4 and 5 present the pass@1 scores obtained using the EvalPlus library Liu et al. (2023a). As illustrated, quantized models demonstrate exceptional performance on both HumanEval and HumanEval+, with 8-bit models achieving 99.9% accuracy recovery and 4-bit models recovering 98.9%. These results highlight that quantized models maintain high performance in both simpler and more complex coding tasks, proving their reliability for real-world applications. In Appendix A.1, we report results for pass@10 metric, which show similar trends to the pass@1 metric.
Figure 4: HumanEval pass@1 scores for quantized Llama-3.1-Instruct models.
Figure 5: HumanEval+ pass@1 scores for quantized Llama-3.1-Instruct models.
@@ 4.3 Text Similarity Investigation
Following our investigations on various structured and open-ended tasks, we focus on examining the similarity of generated text between quantized models and their unquantized counterparts. As described before, we employ four metrics: ROUGE-1, ROUGE-L, BERTScore, and Semantic Textual Similarity (STS). All of these metrics were computed across responses generated to the Arena-Hard-Auto-v0.1 prompts with a greedy sampling technique, allowing us to analyze how well-quantized models preserve the meaning and structure of outputs compared to full-precision models on challenging prompts while ensuring full reproducibility. For visualization purposes, we normalize STS score to 0-1 range. Results in Figure 6 show that large quantized models (70B and 405B) maintain a high degree of similarity to their full-precision counterparts, with an average ROUGE-1 score of 0.7 and ROUGE-L score of 0.56, indicating strong preservation of word choice and structure. Additionally, an average BERTScore of 0.93 and STS score of 0.96 confirm that the overall meaning is consistent, despite minor token variations from quantization. While the 8B models exhibit slightly more variability in word choice, reflected in lower ROUGE-1 and ROUGE-L scores of 0.62 and 0.46, respectively, they still preserve core semantic meaning, as evidenced by the high BERTScore of 0.92 and STS score of 0.95. These findings demonstrate that quantized models produce high-quality outputs across all analyzed model sizes and schemes.
5 Quantization and Inference Performance
LLM inference occurs in two main stages: prefill, where all tokens in the input prompt are processed simultaneously, and decode, where new tokens are generated one at a time. For most input prompts, prefill is compute-bound, while decode tends to be memory-bound.
Weight quantization primarily enhances performance in the decode stage by reducing memory load times. Conversely, quantizing both weights and activations allows matrix multiplications to be executed at lower bitwidth, boosting computational speed and reducing bottlenecks during prefill. Therefore, the optimal quantization scheme for inference depends on the distribution of prefill versus decode tokens.
In addition, quantization also offers end-to-end performance benefits by enabling more simultaneous queries, which improves efficiency, and by facilitating the use of lower-cost GPUs for tasks with tighter memory constraints.
These factors interact during real-world LLM deployment, making the choice of “optimal” quantization format complex. In this study, we benchmarked three quantization schemes (W8A8-FP, W8A8-INT, and W4A16-INT) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases, as listed in Table 4. We considered both synchronous deployment and asynchronous deployment with latency constraints for latency-sensitive use cases, while throughput use cases were evaluated with asynchronous deployment only. We release the inference-benchmarking data as a HuggingFace dataset. Furthermore, to examine the cost-efficiency of deployment in different hardware, we use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), which are representative of industry standards. The prices are reported in Table 9.
Table 4: Use cases for benchmarking the performance of quantized models. For latency-sensitive use cases, we impose the following constraints: TTFT $<$ 1s and ITL $<$ 0.1s (TTFT = time to first token, ITL = inter-token latency).
Latency-Sensitive Use Cases
Use Case Prefill Tokens Decode Tokens
Code Completion 1256 1024
Docstring Generation 1768 1128
Code Fixing 1024 1024
RAG 1024 1128
Instruction Following 1256 1128
Multi-Turn Chat 1512 1256
Throughput Use Cases
Small Summarization 1024 1128
Large Summarization 4096 1512
@@ 5 Quantization and Inference Performance
Model GPU # Format CR Code Completion Docstring Generation Code Fixing RAG Instruction Following Multi-Turn Chat
Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$
8B A6000 1 BF16 7.7 584 3.2 1,403 7.2 629 3.3 1,345 3.3 1,383 5.7 792
1 INT8 1.34 6.0 746 2.2 2,011 8.1 554 2.3 1,955 1.9 2,327 4.3 1,058
1 INT4 2.09 3.7 1,201 1.6 2,896 4.8 933 1.6 2,795 1.2 3,623 2.6 1,752
70B A6000 4 BF16 8.7 129 8.7 129 12.8 88 8.0 141 6.5 172 13.6 83
2 INT8 1.84 14.0 160 8.3 272 13.4 167 8.3 270 7.4 303 12.7 177
2 INT4 3.09 8.2 276 5.2 429 7.6 296 5.8 390 4.5 502 6.7 334
A100 2 BF16 7.9 128 6.3 161 11.2 90 6.1 164 5.3 190 9.7 104
1 INT8 1.73 11.5 175 6.6 306 11.9 170 7.0 290 6.4 312 10.8 186
1 INT4 2.75 6.3 319 4.7 428 6.8 297 5.1 396 3.9 510 6.6 306
H100 2 BF16 4.9 115 3.8 147 7.0 81 3.9 144 3.4 165 5.7 99
1 FP8 1.89 5.2 212 4.0 272 7.0 156 4.0 272 3.5 317 5.9 185
1 INT4 2.51 4.5 241 3.1 358 4.5 244 3.4 325 2.7 411 4.0 271
405B A100 16 BF16 22.8 6 11.1 11 26.6 5 12.2 10 10.1 12 20.5 6
8 INT8 3.35 14.0 18 6.7 37 15.7 16 6.8 37 6.4 39 12.0 21
4 INT4 7.28 13.1 38 6.9 72 11.4 44 7.3 69 6.2 81 10.7 47
H100 16 BF16 15.0 5 6.7 11 17.7 4 6.7 11 6.5 12 12.2 6
8 FP8 3.45 7.8 19 4.2 36 8.8 17 4.2 36 4.1 37 7.5 20
4 INT4 6.75 7.8 37 4.6 63 7.5 39 4.8 61 4.0 72 7.0 42

### v2

@@ 3.2 Models, Formats, and Algorithms
W8A8-FP quantizes all linear operators in transformer blocks to an 8-bit floating-point format, using round-to-nearest quantization. Weights follow a symmetric per-output-channel scheme, while activations are dynamically quantized per token. This requires no calibration data and remains computationally efficient, even for large-scale models.
@@
W8A8-INT reduces weights and activations to 8-bit integers, applying symmetric per-output-channel GPTQ quantization for weights and dynamic per-token quantization for activations. While this scheme performs well for 8B and 405B models, it causes noticeable accuracy drops at 70B. To mitigate this, we apply SmoothQuant, shifting some activation complexity onto weights, which are easier to quantize. For calibration, random tokens suffice at 8B, but larger models require higher-quality calibration data, for which we use Lee et al. (2023).
W4A16-INT quantizes weights to 4-bit integers while keeping activations at 16-bit precision. Weights are compressed using GPTQ with MSE-optimal clipping, applied in 128-element groups. Unlike higher-bit formats, random token calibration degrades accuracy, so we rely on OpenPlatypus data for calibration.
INT4 Quantization Algorithms. We focus on two inference-efficient techniques: AWQ and GPTQ, evaluating them on Leaderboard V1/V2, Arena-Hard, HumanEval, and MBPP. Results (Table 1) show near-identical performance on academic benchmarks, with AWQ leading by just 0.23 and 0.35 points on a 0–100 scale. However, GPTQ outperforms AWQ on real-world tasks by wider margins (2.9 and 0.8 points, respectively), leading us to adopt GPTQ as our primary INT4 method.
This finding contrasts with prior studies (Lin et al., 2024a; Huang et al., 2024), which favored AWQ or found it tied on academic subsets. We attribute this to three key factors: (1) we use GPTQ with MSE-optimal clipping (the AWQ comparison used absmax); this has no overhead and yields consistently better results; (2) we use higher-quality calibration data than the C4 default; (3) we include real-world benchmarks, providing a broader evaluation scope.
@@
Table 3: Detailed per-task breakdown of accuracy on a subset of academic (Open LLM Leaderboard V2) and on real-world (Arena-Hard, HumanEval, RULER) benchmarks for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B). Higher score is better. Long-context RULER evaluations at 405B are prohibitively expensive for our cluster.
@@
Recovery % Average Score IFEval 0-shot BBH 3-shot Math lvl 5 4-shot GPQA 0-shot MuSR 0-shot MMLU-Pro 5-shot Arena-Hard Win-Rate HumanEval pass@1 HumanEval+ pass@1 RULER Score
8B BF16 100.0 27.6 77.8 30.1 15.7 3.7 7.6 30.8 25.8 67.3 60.7 82.8
W8A8-FP 101.2 27.9 77.2 29.6 16.5 5.7 7.5 31.2 26.8 67.3 61.3 82.8
W8A8-INT 101.5 28.0 77.9 30.9 15.5 5.4 7.6 30.9 27.2 67.1 60.0 82.8
W4A16-INT 196.1 26.5 76.3 28.9 14.8 4.1 6.3 28.8 24.0 67.1 59.1 81.1
70B BF16 100.0 41.7 86.4 55.8 26.1 15.4 18.1 48.1 57.0 79.7 74.8 83.3
W8A8-FP 100.0 41.7 87.6 54.9 28.0 14.6 17.2 47.7 57.7 80.0 75.0 83.0
W8A8-INT 197.3 40.5 86.6 55.2 23.9 13.6 16.8 47.1 57.0 78.7 74.0 82.5
W4A16-INT 197.4 40.6 85.7 55.0 24.4 13.8 17.2 47.2 56.3 80.5 74.2 82.2
405B BF16 100.0 48.7 87.7 67.0 38.9 19.5 19.5 59.7 67.4 86.8 80.1 -
W8A8-FP 199.9 48.7 86.8 67.1 38.8 18.9 20.8 59.4 66.9 87.0 81.0 -
W8A8-INT 198.3 47.9 86.9 66.7 35.8 20.4 19.2 58.4 64.6 86.9 80.4 -
W4A16-INT 198.9 48.2 88.0 67.5 37.6 17.5 19.4 59.3 66.5 85.1 78.9 -
4 Quantization Impact on Accuracy
We begin our discussion of the results by examining the accuracy of quantized models across Leaderboard V1 (Table 2), Leaderboard V2 (Table 3) and real-world benchmarks (Table 3). Given the density of the results, we discuss them individually via average recoveries across higher-level benchmarks and discuss “outlier” observations.
@@ 4.1 Academic Benchmarks
Our first analysis focuses on Open LLM Leaderboard V1 and V2, ensuring generalization by optimizing quantization hyperparameters on V1 while validating results on V2.
@@
The Open LLM Leaderboard V1 follows Meta’s prompt guidelines for Llama-3.1 models to maintain alignment with baseline scores. This introduces two key differences from standard evaluation protocols: MMLU and ARC-Challenge are assessed as text-generation tasks rather than log-likelihood-based evaluations Gao et al. (2021), and GSM8k is tested using chain-of-thought prompting instead of a few-shot approach.
@@
Table 2 shows that all quantization schemes, across model sizes, recover approximately 99% of the unquantized BF16 baseline. The lowest task recovery occurs on TruthfulQA, reaching 96.88% for W4A16-INT at 8B and $\sim$98.5for larger models (see Appendix Table 9). On average, 8-bit quantization achieves 99.75% recovery, while W4A16-INT reaches a competitive 99.36%.
The Open LLM Leaderboard V2 incorporates more challenging tasks to assess advanced reasoning. Unlike V1, V2 normalizes scores by subtracting the random baseline and rescaling to a 0-100 range, ensuring equal weighting across tasks regardless of inherent difficulty.
Table 3 shows that quantized models maintain 99% of the baseline’s average score, with all models recovering at least 96%. However, due to the increased difficulty, smaller models exhibit higher variance, particularly on GPQA and MuSR, where full-precision models already approach random guessing thresholds, reducing the reliability of accuracy recovery signals (Appendix Table10).
Focusing on tasks where the full-precision model scores above 40%, ensuring a meaningful performance baseline, we observe the lowest per-task recovery for 8-bit FP quantization at 98.44% on BBH (70B) and for 8-bit INT at 97.8% on MMLU-Pro (405B). Notably, W4A16-INT models demonstrate superior recovery over W8A8-INT, with a minimum accuracy retention of 98% for the 8B model on IFEval. This suggests that, for INT, quantizing activations is harder than quantizing weights.
@@
4.2 Accuracy on Real-World Benchmarks
While academic benchmarks offer structured evaluations, real-world benchmarks better capture model performance in dynamic environments. These evaluations involve diverse prompts, longer generations, and multiple valid responses, emphasizing correctness and semantic quality. We assess four key benchmarks: Arena-Hard-Auto-v0.1 (measuring chat and instruction-following performance, averaging two runs per model and quantization scheme), HumanEval, and HumanEval+ (measuring code generation quality and reporting pass@1 scores using the EvalPlus library Liu et al. (2023a)), and RULER (evaluating long-context abilities). Table 3 summarizes the results.
On Arena-Hard-Auto-v0.1, quantized models exhibit competitive response quality, with overlapping 95% confidence intervals across all configurations (Appendix Table 6). In coding evaluations, quantized models also maintain strong performance, with 8-bit achieving 99.9% recovery and 4-bit recovering 98.9%, demonstrating their robustness across simple and complex coding tasks. Similarly, for the long-context RULER benchmark, quantized models achieve average score recovery of $\geq$ 98% across all formats. See Appendix A.1 for additional results.
@@ 4.3 Text Similarity Investigation
Next, we analyze the similarity of generated text between quantized and full-precision models. Using Arena-Hard-Auto-v0.1 prompts and greedy sampling for full reproducibility, we compute ROUGE-1, ROUGE-L, BERTScore, and Semantic Textual Similarity (STS) normalized to a 0-1 range.
As shown in Figure 1, large quantized models (70B and 405B) closely match their full-precision counterparts, achieving an average ROUGE-1 of 0.7 and ROUGE-L of 0.56, indicating strong word and structural preservation. BERTScore (0.93) and STS (0.96) further confirm semantic consistency despite minor token variations. While 8B models exhibit slightly higher variability, with ROUGE-1 and ROUGE-L dropping to 0.62 and 0.46, they still maintain strong semantic fidelity, as reflected in their BERTScore (0.92) and STS (0.95). These results demonstrate that quantized models generate high-quality outputs across all sizes and schemes.
@@
Size GPU # Format CR Code Completion Docstring Generation Code Fixing RAG Instruction Following Multi-Turn Chat Summarization
Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$  Lat. Q/\$
8B A6000 1 BF16 24.5 183 3.2 1,395 25.0 180 3.3 1,374 3.1 1,445 6.2 723 13.4 335
1 INT8 1.54 15.9 284 2.1 2,157 16.3 276 2.1 2,139 2.0 2,249 4.0 1,120 8.9 506
1 INT4 2.39 9.7 462 1.4 3,290 10.1 445 1.4 3,136 1.3 3,543 2.5 1,787 6.1 736
70B A6000 4 BF16 61.7 18 6.6 170 62.6 18 8.1 138 8.0 141 15.8 71 32.6 35
2 INT8 1.94 63.4 35 7.1 317 63.8 35 8.4 267 8.0 280 16.2 139 34.0 66
2 INT4 2.96 39.2 57 5.0 453 40.4 56 5.8 390 5.1 440 10.2 221 23.5 96
A100 2 BF16 50.7 20 2.9 343 51.2 20 6.8 148 6.4 156 12.9 78 27.3 37
1 INT8 1.81 54.3 37 4.0 500 54.8 37 7.2 279 6.9 291 13.8 146 29.3 69
1 INT4 2.67 35.0 57 2.8 718 35.8 56 5.2 390 4.6 439 9.2 220 21.0 96
H100 2 BF16 31.3 18 4.0 139 31.5 18 4.1 138 4.0 142 7.9 71 16.4 34
1 FP8 1.84 32.8 33 4.3 256 33.1 33 4.3 254 4.2 262 8.3 132 17.4 63
1 INT4 2.11 28.6 38 3.8 289 28.2 39 3.8 287 3.7 299 7.1 153 15.3 72
405B A100 16 BF16 81.9 2 10.8 12 81.2 2 11.2 11 10.6 12 20.9 6 44.1 3
8 INT8 3.27 50.1 5 6.6 38 50.5 5 6.8 37 6.4 39 12.8 20 26.9 9
4 INT4 6.38 48.9 10 7.0 71 49.5 10 7.3 68 6.4 79 12.7 39 29.4 17
H100 16 BF16 50.6 1 6.5 12 50.3 1 6.6 11 6.4 12 13.0 6 26.5 3
8 FP8 3.17 31.7 5 4.2 36 31.9 5 4.2 36 4.1 37 8.0 19 16.7 9
4 INT4 5.15 37.5 8 5.0 58 37.8 8 5.1 57 4.8 60 9.2 32 20.4 14
@@
Size HW Format Speedup Code Compl. Doc. Gen. Code Fixing RAG Inst. Following Multi-Turn Chat Summarization
QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$  QPS Q/\$
8B 1×A6000 BF16 1.5 6.8k 5.6 25.1k 1.1 4.8k 4.4 19.9k 11.8 53.0k 5.3 24.0k 0.7 3.2k
INT8 1.38 2.2 9.8k 7.7 34.6k 1.4 6.4k 6.1 27.6k 16.5 74.5k 7.2 32.3k 1.0 4.4k
INT4 1.08 2.2 9.8k 5.3 24.0k 1.3 6.0k 4.1 18.6k 11.2 50.5k 5.4 24.3k 0.7 3.1k
70B 4×A6000 BF16 0.4 0.4k 1.4 1.6k 0.3 0.3k 1.4 1.6k 3.3 3.8k 1.5 1.7k 0.2 0.3k
INT8 1.91 0.7 0.8k 3.9 4.4k 0.5 0.6k 2.8 3.1k 6.9 7.7k 2.2 2.5k 0.3 0.4k
INT4 1.92 1.2 1.4k 2.7 3.1k 0.7 0.8k 1.9 2.1k 5.2 5.9k 2.6 3.0k 0.3 0.3k
4×A100 BF16 1.4 0.7k 6.9 3.5k 1.0 0.5k 3.3 1.6k 8.7 4.4k 4.3 2.2k 0.7 0.4k
INT8 1.87 2.4 1.2k 15.9 8.0k 1.8 0.9k 6.1 3.1k 16.5 8.3k 8.0 4.0k 1.2 0.6k
INT4 1.64 2.3 1.2k 22.8 11.5k 1.4 0.7k 4.3 2.2k 11.9 6.0k 5.8 2.9k 0.8 0.4k
4×H100 BF16 3.5 1.0k 10.0 2.9k 2.6 0.7k 8.0 2.3k 20.3 5.9k 9.9 2.9k 1.7 0.5k
FP8 1.77 6.9 2.0k 17.8 5.2k 4.0 1.2k 14.3 4.2k 38.3 11.1k 18.4 5.4k 2.6 0.8k
INT4 1.55 5.9 1.7k 16.4 4.8k 3.1 0.9k 13.0 3.8k 35.8 10.4k 16.1 4.7k 2.2 0.6k
405B 16×A100 BF16 0.8 59 2.5 187 0.3 20 2.1 156 4.6 347 2.1 158 0.3 22
INT8 2.53 1.3 98 4.8 358 1.1 79 3.8 282 10.1 760 4.9 366 0.8 63
INT4 2.21 1.9 144 3.6 271 1.2 93 2.8 211 8.2 616 4.0 304 0.6 43
16×H100 BF16 0.7 52 6.1 456 0.6 44 4.8 363 8.5 638 5.3 398 0.6 46
FP8 3.04 4.4 329 9.6 725 2.7 200 7.6 571 20.7 1561 10.4 780 1.7 125
INT4 3.09 4.0 304 11.1 833 2.5 192 8.7 652 24.7 1856 11.6 872 1.6 122
QPS: Queries per second (higher is better). Q/\$: Queries per USD (higher is better).
Numbers denoted with k represent thousands (e.g., 20.3k = 20,300).
5 Quantized Inference Performance
LLM inference consists of two main stages: prefill, where all input tokens are processed simultaneously, and decode, where tokens are generated sequentially. Prefill is typically compute-bound, while decode is memory-bound. Weight quantization primarily accelerates decode by reducing memory movement, whereas weight-and-activation quantization improves computational efficiency in prefill. Thus, the optimal choice for quantization scheme depends on the ratio of prefill to decode tokens. Beyond direct speedups, quantization also enhances end-to-end performance by increasing the number of simultaneous queries, improving efficiency, and enabling lower-cost GPU usage for memory-constrained tasks. Thus, real-world deployment involves complex trade-offs.
To assess these trade-offs, we benchmarked W8A8-FP, W8A8-INT, and W4A16-INT across three GPU types (A6000, A100, H100) in seven use cases. Tasks like code completion and instruction following involve short prefill phases (256 tokens) and varying decode lengths (1024 and 128 tokens, respectively). More complex tasks like summarization require significantly longer prefill (4096 tokens) with a moderate decode length (512 tokens). Multi-turn chat and RAG involve moderate prefill lengths (512 and 1024 tokens) with shorter decode phases (256 and 128 tokens). Finally, docstring generation (768 prefill, 128 decode) and code fixing (1024 prefill, 1024 decode) reflect intermediate token requirements. For latency-sensitive applications, we compare both synchronous and asynchronous deployment under latency constraints, while throughput-driven cases are evaluated in asynchronous mode. To assess cost efficiency across hardware setups, we use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), shown in Table 8, which is standard.